A 200-kg piano on frictionless casters is to be pushed up a 7-m ramp to a loading platform 1 m above the ground. How much force is needed? 


1
Expert's answer
2021-11-03T18:36:30-0400

"F=mg\sin\alpha=200\cdot9.81\cdot(1/7)=280.3\ (N)"